OpenStudy (anonymous):

1. Solve 5x² = -45x.

OpenStudy (anonymous):

x=0 or x=-9

OpenStudy (anonymous):

5x (x + 9) = 0

OpenStudy (anonymous):

How do yo work it out?

OpenStudy (anonymous):

either 5x = 0 or x+9 = 0

OpenStudy (anonymous):

so x = 0 or -9

OpenStudy (anonymous):

\[5x^2=-45x\] divide both sides by 5 \[x^2=-9x\] add 9x to both sides \[x^2+9x=0\] factor \[x(x+9)=0\] set each factor equal to zero and solve \[x=0,x+9=0\implies x=-9\]

OpenStudy (anonymous):

Thank you!!